7 green 8 blue one white marble
find the probabilty of 2 blue marbles in a row without replacement

Answer by checkley77(12844)

8/16*7/15=1/2*7/15=7/30=.2333=23.33 % IS THE PROBABILITY OF 2 BLUES IN A ROW.